OpenStudy (vortish):

solve the system of equations by the substitution method x + y = 12 y = -5x

OpenStudy (anonymous):

x+y=12 y=-5x > x-5x=12 -4x=12 x=-3 > y-3=12 y=15

OpenStudy (mathstudent55):

Since the second equation is arleady solved for y, just insert what y is equal to into the first equation: x + y = 12 x - 5x = 12 Then solve for x. Then insert what x equals into either the first or second equation, and solve for y.

OpenStudy (vortish):

the question has a multiple answer thing and those are A) the solution is ( ) in ordered pair B) there are infantly many solutions C0 there is no solution

OpenStudy (vortish):

so if y = 15 would the ordered pair be (15,-3)

OpenStudy (anonymous):

(x,y) so, (-3,15)